OpenStudy (jenessa_2413):

During his workouts last week, Allen jogged the same amount of time each day for 6 days. Use the following information to determine the number of minutes Allen jogged each day: During each workout, he jogged and he rode his bike. He rode his bike for 45 minutes each day. He worked out for a total of 360 minutes last week. What is the number of minutes that Allen jogged each day last week? Enter your answer in the box below.

OpenStudy (welshfella):

how many minutes did he spend on his bike in the 6 days?

OpenStudy (welshfella):

he biked 45 minutes each day so in 6 days he road the bike for how many minutes?

OpenStudy (jenessa_2413):

is it 15 minutes?

OpenStudy (jenessa_2413):

@welshfella

OpenStudy (welshfella):

how did you get that - is that the final answer?

OpenStudy (welshfella):

How many minutes did he spend working out each day?

OpenStudy (jenessa_2413):

90 minutes of jogging divided by 6 days = 15 minutes He jogged 15 minutes each day. is this right?

OpenStudy (jenessa_2413):

@welshfella

OpenStudy (welshfella):

yes

OpenStudy (jenessa_2413):

ok thanks

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

360 - 45*6 ---------- = (360 - 270) / 6 = 90 / 6 = 15 6